OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the graph of the equation 2x - y = 8. intercept = (2, 0); y-intercept = (0, -8) x-intercept = (4, 0); y-intercept = (0, -8) intercept = (2, 0); y-intercept = (0, 8) intercept = (4, 0); y-intercept = (0, 8)

OpenStudy (anonymous):

First you should solve the equation if you haven't already...

OpenStudy (anonymous):

will you walk me through it? my teacher didnt help me at all

OpenStudy (anonymous):

No problem. 2x-y=8 (subtract 2x from both sides) -y=-2x+8 (multiply both sides by -1) y=2x-8 At this point you have your y-intercept which is (0,-8)

OpenStudy (anonymous):

okay thank you like i have needed alot o help on math

OpenStudy (anonymous):

Me too. Do you know how to find the x-intercept from here?

OpenStudy (anonymous):

add 1 on both sides?

OpenStudy (anonymous):

You have to substitute y for 0 so it should look like 0=2x-8 then solve for x like you solved for y before.

OpenStudy (anonymous):

x=4?

OpenStudy (anonymous):

so b?

OpenStudy (anonymous):

Yep.